OpenStudy (anonymous):

A box that hold a pen has square ends, and its length is 13cm longer than its width. The volume of the box is 60cm^3. Determine the dimensions of the box. How am i suppose to find the height, or any of the dimensions based on what i am given

OpenStudy (ranga):

The ends are square. So the width and the height are the same. Assume the width is W Length L = (W + 13) Volume = Length x Width x Height = (W + 13)(W)(W) = 60 Solve for W. Then find L.

OpenStudy (anonymous):

How do i solve for w?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Multiply and expand the equation: W^2(W + 13) = 60 W^3 + 13W^2 - 60 = 0 This is a cubic equation. It has to be solved by trial and error. Try W = 1, W = 2, W = 3, etc. and see when W^3 + 13W^2 - 60 changes sign. As soon as you see the sign change you will know there is a root (or a zero) somewhere between them. Can you tell me what you get for (W^3 + 13W^2 - 60) when you try W = 1, W = 2, W = 3?

OpenStudy (anonymous):

Thank you so much! This totally helped me! It will help me for all my other questions too! Thank you again!

OpenStudy (ranga):

You are welcome. So what did you get for W and L?

OpenStudy (anonymous):

the width and height are 2cm, and the length is 15cm!

OpenStudy (ranga):

Yes, you have the correct answer! Luckily here for W = 2, (W^3 + 13W^2 - 60) goes to zero. So we know W = 2 is the solution. If the expression did not go to 0 then we have to see for between what values of W the expression changes sign (goes from positive to negative or negative to positive) and then try decimal values for W in between those points. It is good we did not have to do that here.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.